OpenStudy (anonymous):

Which choice could be represented by the expression mc005-1.jpg You buy a present for $48, leaving you with $336. How much money did you have to begin with? You and your friends are evenly dividing $336. Each of you gets $48. How many friends are you splitting the money with? You need to save $48 more dollars before you have enough money to buy the computer you want for $336. How much do you have saved?

OpenStudy (anonymous):

mc005-1.jdg= x+48=336